# Universal Omni-Language Models

### v3

@@ 5.1 Results on Omni-Language Models
We present the breakdown of the performance of open-source omni-language model baselines across different audio types and task categories in the OmniBench evaluation. The results reveal inconsistent performance patterns across audio types, with some models showing higher accuracy on sound events or music compared to speech. Across task categories, models tend to perform better on object identification and description tasks, while struggling with more reasoning tasks such as plot inference and story description, as illustrated by Table 3. This might be because visual entity recognition is an essential component for image captioning and other type of pre-training dataset. We observe Gemini provides significantly better results on the context/environment entities other than object entities. Furthermore, most of the models perform really bad on quantity & counting tasks. But scaling up of UnifiedIO model contributes a lot to this type of task. And scaling up from 1.1B to 3.2B benefits all senarios.
These findings highlight current limitations of OLMs in integrating information across modalities.
5.2 Textual Approximation on Images and Audios
@@
As the absence of strong OLM baselines on OmniBench, we further introduce the text alternatives of images ($I^{\prime}$) and audios ($A^{\prime}$) to embrace more dual-modal MLLMs to analyze the current research progress in the field. The results using audio transcripts and image captions are put in Table 4, Table 5 correspondingly (results of $(I^{\prime},A^{\prime})$ setting at Table 6, Appendix A).
@@
In the setting of using text as the alternatives of audios and images, the VLMs show generally better results than ALMs (Table 4 vs Table 5) even compared with open-source model with similar model size. This could be caused by : 1) more available research resources have been put in VLMs to develop datasets and cross-modality alignment architectures, leading to higher instruction following rate and accuracy compared to ALMs; 2) the audio data are naturally harder (and hence more expensive) to annotate; and 3) audio typically has longer sequence tokens and requires more computational resource compared to text and image, making it harder to scale up. If $I^{\prime}$ and $A^{\prime}$ have the information loss ratio when converted from $I$ and $A$, it seems to be easier for the researchers to train the future omni-language models from exisiting VLMs rather than ALMs. Besides, we can observe Claude-3.5 and GPT-4o are generally the best two VLMs, significantly better compared to open-source VLMs. And Qwen2-audio and Gemini are the best two ALMs in speech and audio, and Audio-SALMONN is the best on music. Moreover, we can see significant difference on different type of audio, i.e., LTU and audio-flamingo are worse for music compare to speech and audio, while Qwen-audio which include music on pre-training provides better results on music compared to speech. And MusiLingo only use music for pre-training perform worse in speech and sudio.
@@
Pure Textual Evaluation.
The performance gaps brought by the replaced textual image and audio descriptions are in revealed in Figure 5. Notably, the majority of models demonstrate improved accuracy when processing textual representations of multimodal data compared to their performance on either image captions or audio transcripts alone. This suggests that these models show stronger reasoning capability when equipped with information from multiple textual sources rather than handling raw multimodal inputs. For instance, Qwen2-Audio-7B-Instruct shows a significant jump in accuracy from 39.05% (audio transcript only) and 39.67% (image caption only) to 47.02% when given both textual representations. Similarly, proprietary models like GPT4-o and Claude-3.5-Sonnet exhibit substantial gains, with GPT4-o (0513) achieving an impressive 60.60% accuracy in the pure textual setting.
@@ 6 Conclusion and Future Study
Ethics Statement
Our research on OmniBench and the development of multimodal language models raises several important ethical considerations:
  • Data Collection and Privacy: All image and audio data used in OmniBench was collected from public sources or created specifically for this research. We took care to remove any personally identifiable information. For human-recorded audio, participants provided informed consent and were compensated fairly for their time.
  • Potential Biases: We acknowledge that the dataset may contain inherent biases in terms of language, cultural representation, and types of scenarios depicted. We have made efforts to include diverse content, but further work is needed to fully characterize and mitigate these biases. Users of OmniBench should be aware of these limitations.
  • Responsible Disclosure: We will release OmniBench publicly to foster open research, but with appropriate use guidelines. The OmniInstruct dataset will be made available to researchers who agree to terms of responsible use.
We are committed to ongoing evaluation of the ethical implications of this work as the field of multimodal AI continues to advance rapidly.
Reproducibility Statement
We have made significant efforts to ensure the reproducibility of our work on OmniBench and the associated experiments:
  • Dataset: The complete OmniBench dataset, including all images, audio files, and question-answer pairs, will be made publicly available upon publication. Detailed information about the data collection process, annotation guidelines, and quality control measures are provided in section 3 and Appendix B.
  • Code: We have developed and will release a comprehensive codebase that includes: Scripts for data preprocessing and formatting; Implementation of all evaluation metrics; Code for running experiments.
  • Model Evaluation: For all baseline models evaluated, we provide detailed specifications. For proprietary models, we specify the exact API versions used and the dates of access.
  • Reproducibility Challenges: We acknowledge that exact reproduction of results for some proprietary models may be challenging due to potential API changes.
By providing these resources and detailed documentation, we aim to facilitate the reproduction of our results and encourage further research in this area. We welcome feedback from the community on any aspects that require additional clarification to ensure full reproducibility.
@@ B.1 Statistics for OmniInstruct Dataset
MSRVTT-QA 140,554 11,143 80,078 (57.0%) 6,479 (58.1%)
Total 233,206 27,941 84,580 8,390
As demonstrate in Table 7, most of the samples in the dataset are in low quality and therefore abandoned, and only 93k of samples remain for Omni-modality SFT training. This is reasonable because most of the questions are generated from templates, and the image may not sampled from the most relevant part of the questions and, therefore hot high in quality.

### v4

@@ 5.1 Results on Omni-Language Models
We present the breakdown of the performance of open-source omni-language model baselines across different audio types and task categories in the OmniBench evaluation in Table 4 and Table 5. The results reveal inconsistent performance patterns across audio types. For instance, despite overall poor performance, open-source baselines generally exhibit higher accuracy on speech audio, indicating a potential bias towards speech data. Besides, Video-Salmonn and Gemini-1.5-Pro provide better results on music subsets compared to their performance on speech and music, potentially due to their large corpus of music videos, though the music ethics of training foundation models are still under-discussion (Ma et al., 2024). Across task categories, many of the models, such as Gemini-1.5-Pro, tend to perform better on object identification and description tasks while struggling with more reasoning tasks such as plot inference and story description. This might be because visual entity recognition task is an essential component for image captioning and other type of pre-training dataset. Furthermore, some models like Gemini-1.5-Pro, Reka-core-20240501, and Video-SALMONN perform significantly badly on quantity & counting tasks compared with the rest of the tasks. But scaling up of UnifiedIO model contributes a lot to this type of task.
Human Evaluation
We invite three human annotators to evaluate on the OmniBench questions, deriving a much higher accuracy $63.19\%$ compared to all the OLMs333 Human performance is limited by time constraints, annotators’ non-native language or music knowledge gaps, and their associate-level education. . The human evaluator results hold a Fleiss’ Kappa value of $0.421$ suggests, suggesting a high level of inter-annotator agreement. Different from models, human evaluation shows a particularly good performance on ”Sound Event” among the audio types, which is reasonable since the sounds are short and straightforward without requiring complex reasoning to recognize. Moreover, human shows higher scores on ”Abstract Concept” tasks, potentially because these tasks require more reasoning.
5.2 The Effectiveness of OmniInstruct
To validate the effectiveness of OmniInstruct, we conducted experiments using MIO-instruct, a 7B parameter omni-language model previously trained only on audio-language and vision-language pairs. We evaluated the impact of supervised fine-tuning using OmniInstruct through two experimental settings. First, we created a compact subset of OmniInstruct containing 6.4K samples (approximately 7.5% of the full dataset) to enable efficient experimentation. In our first setting (MIO-Instruct-Omni-V1), we directly fine-tuned the model using its vanilla multimodal tokenizers. In the second setting (MIO-Instruct-Omni-V1-voice-filtered), we refined the approach by filtering out non-speech tokens from the audio RVQ tokenizer that only supports speech, aligning with MIO-Instruct’s original training to avoid mode collapse. The results demonstrate meaningful improvements: while the baseline MIO-Instruct achieved 24.8% accuracy on OmniBench, fine-tuning with the vanilla setting improved performance to 25.7%, and the voice-filtered approach further enhanced accuracy to 29.2%. These results indicate that even a small subset of OmniInstruct can effectively adapt existing multimodal language models to tri-modal scenarios. Notably, the superior performance of the voice-filtered approach suggests that aligning fine-tuning data distribution with a model’s distribution can yield better results.
@@ 5.2 The Effectiveness of OmniInstruct
5.3 Textual Approximation on Images and Audios
As the absence of strong OLM baselines, we further introduce the text alternatives of images ($I^{\prime}$) and audios ($A^{\prime}$) to embrace more dual-modal MLLMs to analyze the current research progress in the field.
@@ 5.3 Textual Approximation on Images and Audios
In the setting of using text as the alternatives of audios and images, the VLMs show generally better results than ALMs (Table 6 vs Table 7) even compared with open-source model with similar model size [^4] The results of pure text $(I^{\prime},A^{\prime})$ setting are placed at Table 8, Appendix A.. This could be caused by : 1) more available research resources have been put in VLMs to develop datasets and cross-modality alignment architectures, leading to higher instruction following rate and accuracy compared to ALMs; 2) the audio data are naturally harder (and hence more expensive) to annotate; and 3) audio typically has longer sequence tokens and requires more computational resource compared to text and image, making it harder to scale up. If $I^{\prime}$ and $A^{\prime}$ have the information loss ratio when converted from $I$ and $A$, it seems to be easier for the researchers to train the future omni-language models from exisiting VLMs rather than ALMs. Besides, we can observe Claude-3.5 and GPT-4o are generally the best two VLMs, significantly better compared to open-source VLMs. Qwen2- audio and Gemini are the two best ALMs in speech, Qwen2- audio is the best in sound, and Audio-SALMONN is music. Moreover, we can see significant differences in different types of audio, i.e., LTU and audio-flamingo are worse for music compared to speech, while Qwen-audio, which includes music on pre-training, provides better results on music compared to speech. And MusiLingo only uses music for pre-training performs worse in speech and audio.
Pure Textual Evaluation.
The performance gaps brought by the replaced textual image and audio descriptions are in revealed in Figure 4. Notably, the majority of models demonstrate improved accuracy when processing textual representations of multimodal data compared to their performance on either image captions or audio transcripts alone. This suggests that these models show stronger reasoning capability when equipped with information from multiple textual sources rather than handling raw multimodal inputs. For instance, Qwen2-Audio-7B-Instruct shows a significant jump in accuracy from 39.05% (audio transcript only) and 39.67% (image caption only) to 47.02% when given both textual representations. But the gains of best proprietary models GPT4-o (0513) andClaude-3.5-Sonnet exhibit is not significant, though GPT4-o (0513) achieved an impressive 60.60% accuracy in the pure textual setting.
@@ 6 Conclusion and Future Study
Impact Statement
This paper introduces OmniBench, a crucial step towards developing truly multimodal AI. By rigorously evaluating models on their ability to integrate visual, acoustic, and textual information, OmniBench exposes critical limitations in current approaches and highlights the need for dedicated research in tri-modal reasoning. This benchmark has significant potential societal impact, as robust OLMs could revolutionize fields like healthcare (improved diagnostics), accessibility (enhanced assistive technologies), and human-computer interaction (more natural interfaces). Furthermore, the accompanying OmniInstruct dataset provides a valuable resource for training and fine-tuning these models. While the development of OLMs raises ethical considerations regarding bias and fairness, OmniBench’s focus on comprehensive evaluation enables researchers to identify and address these issues proactively. By fostering transparency and reproducibility, this work paves the way for responsible development and deployment of powerful multimodal AI systems that benefit society.
@@ B.1 Statistics for OmniInstruct Dataset
MSRVTT-QA 140,554 11,143 80,078 (57.0%) 9,614 (86.2%)
Total 233,206 27,941 84,580 11,525
As demonstrate in Table 9, most of the samples in the dataset are in low quality and therefore abandoned, and only 96k of samples remain for Omni-modality SFT training. This is reasonable because most of the questions are generated from templates, and the image may not sampled from the most relevant part of the questions and, therefore hot high in quality.
@@ B.3 Diversity of Music Audios of OmniBenchmark
Appendix C Statistic Significance of the Findings
In this paper, we provide about 30 claims on the discussion in section 5. In order to discover whether the observation is due to the bias of a small test set, We utilize multiple types of hypothesis testing for these experimental results.
C.1 Inference if Two Models Provide Significantly Different Results under the Same Small Subset
The method is inspired by  (Guyon et al., 1998). Suppose all the data samples $\{x_{i}\}_{i=1}^{n}$ are i.id, where $n$ is the number of data samples. For each specific model $M_{j}$, the performance of the $j^{th}$ model on the $i^{th}$ data is $M_{j}\left(x_{i}\right)$ which is equal to 00 if the model predicts the wrong choice and $1$ otherwise. Due to the i.id. assumption of the test set, the $\{M_{j}\left(x_{i}\right)\}_{i=1}^{n}$ is a random variable sequence (r.v.s.) with binomial distribution with mean $p_{i}$ and variance $\sigma_{i}^{2}=p_{i}(1-p_{i})$.
Suppose the accuracy $p_{i}$ of a given model is estimated by computing the average correct rate $\hat{p_{i}}$ over a finite number $n$ of test examples or patterns. When $n$ is small, $\hat{p_{1}}>\hat{p_{2}}$ may hold even when ${p_{1}}<{p_{2}}$.
According to the normal law, the whited random variable
$Z:=\frac{p-\hat{p}}{\sigma\/\sqrt{n}}$
obeys the standardized Normal law (with mean 0 and variance 1). Define the real number $Z_{\alpha}$ to be the value such that the probability $\mathbb{P}\left(x>Z_{\alpha}\right)=\alpha$. Then, according to Chebychev’s inequality,
$\mathbb{P}\left(p-\hat{p}\leq\frac{\sigma z_{alpha}}{\sqrt{n}}\right)\leq\alpha$
We are then calculating the number of test examples $n$ that is needed to guarantee a certain margin of error $\epsilon_{i}$, e.g., $\epsilon(n,\alpha)=\frac{z_{\alpha}\sigma}{\sqrt{n}}$for the Normal law). Denote $\beta:=\frac{\epsilon_{i}}{p_{i}}$. Then
$p-\hat{p}\leq z_{\alpha}\sqrt{\frac{p(1-p)}{n}}$
hold with the probability $1-\alpha$.
Therefore,
$n\leq\left(\frac{z_{\alpha}}{\beta}\right)^{2}\frac{1-p}{p}$
holds with the probability $1-\alpha$.
According to the cdf of standard normal distribution,
$n\frac{\beta^{2}p_{1}}{1-p_{1}}=\frac{n(p_{2}-p_{1})^{2}}{p_{1}(1-p_{1})}\leq 2.72$
implies significance 0.05, and the value $\leq 5.29$ implies p-value less than 0.01.
SImilarly, with Chernoff bound  (Chernoff, 1952), $p-\hat{p}\leq\sqrt{\frac{-2p\left(\ln\alpha\right)}{n}}$ hold with the probability $1-\alpha$. And therefore, $n=\frac{-2\ln\alpha}{\beta^{2}p}$. The alternative estimation of p-value is $\alpha=\exp\left(\frac{np\beta^{2}}{2}\right)$
Therefore, we have the following claim with significance value:
  • In table2
    • Claim (1): Scaling up may not contribute to the overall performance. With audio and image as input, UnifiedIO2-xlarge’s performance 38.00% is better than the scaled-up UnifiedIO2-xxlarge’s (33.98%) version on the 1142 samples holds, with a p-value is less than 0.01.
    • Claim (2): Performance f Gemini-1.5-Pro with image and audio input (42.91%) is better than single modality input (e.g. 27.93% for image only) on the 1142 samples holds with p-value 1.19e-20.
  • Table 3
    • Claim (3): Scaling up of UnifiedIO can contribute to Count & Quantity tasks, though only 15 samples are selected. UnifiedIO-large (6.67%) is worse than UnifiedIO-xlarge (20%) holds with a p-value less than 0.05, and UnifiedIO-xlarge (20%) is worse than UnifiedIO-xxlarge (46.67%) holds with a p-value less than 0.01.
  • In table 4 & table 5
    • - Claim (5): The performance of Claude-3.5 (59.37%) and GPT-4o (57.62%) in Table 4 is better than open-source VLM (less or equal to 54.29%). We can say Claude-3.5 is better than InternVL-2-40B (54.29%) with a p-value less than 0.05, and GPT-4o is better than the second-best open-source VLM InternVL-2-26B (51.75%) with p-value less than 0.01. But we can not say GPT-4o (57.62%) is better than InternVL-2-40B (54.29%).
    • Claim (6): **In speech subset (771 samples)** demonstrated in table 3 with audio and image as input, Gemini-1.5-Pro (42.67%) does not surpass UnifiedIO2-xlarge (3.2B) (39.56%) significantly, but it surpass all other models like the second-best UnifiedIO2-xxlarge (6.8B) (34.24%) with p-value less than 0.01
    • Claim (7): **In general audio subset (265 samples)** demonstrated in Table 3 with audio and image as input, Gemini-1.5-Pro (42.26%) is better than all the others besides UnifiedIO2, such as Video-SALMONN (13B) (31.70%) with p-value less than 0.01
    • Claim (8): **In the music subset (106 samples)** demonstrated in Table 3 with audio and image as input, Video-SALMONN (13B) (56.6%) is better than all the other models such as the best one Gemini-1.5-Pro (46.23%) with p-value less than 0.05.
    • Claim (9): **In the speech subset (771 samples)** demonstrated in Table 5 with audio and image caption as input, Qwen2-audio (40.60%) and Gemini (39.82%) are better than all other ALMs. E.g. Gemini is better than the best of the rest Audio-SALMONN (13B) (34.5%) with p-value less than 0.01
    • Claim (10): **In general audio subset (265 samples )** demonstrated in table 5 with audio and image caption as input, Qwen2-audio (41.89%) is better than every model besides UnifiedIO2-xxlarge (38.87%), e.g. better than Gemini-1.5-Pro (33.96%) with p-value less than 0.05.
    • Claim (11): **In music subset (106 samples)** demonstrated in table 5 with audio and image caption as input, Audio-SALMONN (50%) is better than all the others besides Gemini-1.5-Pro (41.5%). For example, Audio-SALMONN is better than the best of the rest Qwen2-Audio-7B-Instruct (38.68%) with a p-value less than 0.05.
  • Compared with Table 6
    • Claim (12): In Table 6, with audio transcription and image caption as inputs, Qwen2-Audio-7B-Instruct (47.02%) is better than with only audio transcription (39.05%) or only image caption (39.67%) with a p-value less than 0.01
    • Claim (13): Replace the input of Qwen2-Audio-7B-Instruct from image caption and audio in Table 5 (40.72%) to caption and audio transcription in Table 6 (47.02%), the performance increases significantly with a p-value less than 0.01
    • Claim (14): Replace the input of Claude-3.5-Sonnet from image and audio transcription in Table 4 (59.37%) to image caption and transcription in Table 6 (56.83%) has no significant changes. And GPT4-o (0513) performances in Table 4 (57.62%) and Table 6 (60.51%) has no significant changes either.
    • Claim (15): GPT4-o (0513) performance in Table 6 (60.51%) has no significant difference with human (63.19%)
C.2 Comparasion between Two Different Setting with the Same Samples
We use student-paired t-tests to compare two different experimental settings on the same testset.
Recaall the performance of UnifiedIO2 in different experimental settings are as follows:
  • performance with image and audio (25.94%, 39.56%, 34.24% in Table 2)
  • performance with image caption & audio (29.16%, 32.22%, 32.05% in Table 5)
  • performance with image & audio traiscripton (34.33%, 43.17%, 40.81% in Table 4)
  • performance with pure text input (30.74%, 33.80%, 34.15% in Table 6)
Then the t-test results are as follows:
  • between Table 2 and Table 4 with audio transcription has a p-value of 0.0471, showing the model has room for improvement in audio understanding capability compared with audio transcription ground truth.
  • Claim (17): There is no significant difference between Table 2 and Table 5 (p-value 0.562) or Table 6 (p-value 0.919), showing the capability of OLM on image understanding is similar to the SOTA LLM generated image caption.
C.3 Comparasion on the Same Setting with Two Different Subset Samples
We utilize Wilcoxon rank sum test (i.e. Mann–Whitney U test) to compare the model performance on two different subsets with different but independent samples.
  • In table 3 (top):
    • Claim (18): The performance of salmonn on the music subset (56.60%, 106 samples) is better than speech (34.11%, 771 samples) with p-value 6.84e-6; and better than sound events (31.70%, 265 samples) with p-value 8.98e-6.
    • Claim (19): The performance of UnifiedIO2-xlarge (3.2B) music (29.25%) may not be worse than sound (36.98%) with p-value 0.158; but may be worse than speech (39.56%) with p-value 0.0407.
    • Claim (20): The performance of Reka-core-20240501 on sound (26.04%) may not be worse than music (33.02%) with p-value 0.177; and may not be worse than speech (31.52%) with p-value 0.093.
  • table3 (bottom)
    • Claim (21): the performance of Gemini-1.5-Pro on Story Description (30.87%, 230 samples) is worse than all the other 7 class with p-value 5.38e-5; Plot Inference (32.91%, 237 samples) worse than others with p-value 6.69e-4; Object Identification & Description (62.56%, 211 samples) is better than others with p-value 9.46e-11. This statement holds for many other models, but not for AnyGPT.
    • Claim (22): Gemini-1.5-Pro on Count & Quantity (13.33%, 15 samples) is worse than all the other 7 classes with a p-value of 0.0209. Reka-core-20240501 on Count & Quantity (6.67%) is worse than the rest with a p-value of 0.0445. Video-SALMONN on Count & Quantity (6.67%) is worse than the rest with a p-value of 0.0239. But not for AnyGPT.
  • In table 5
    • claim (23): The performance of LTU on music (16.04%) is worse than speech (25.42%) with a p-value of 0.0348, but may not be worse than sound (20.00%) with a p-value of 0.379.
    • claim (24): The performance of audio-flamingo on music (16.98%) is worse than speech (26.98%) with a p-value of 0.0275, but may not be worse than sound (21.51%) with a p-value of 0.328.
    • claim (25): The performance of Qwen-audio on music (25.47%) is better than speech (14.66%) with a p-value of 0.0044.
    • claim (26): The performance of MusiLingo on music (25.47%) is better than speech (11.93%) with a p-value of 1.37e-4, and better than sound (13.96%) with a p-value of 5.96e-3.
Given that we conducted 20-30 hypothesis tests, it is not surprising to observe at least one p-value below 0.05 due to the multiple comparisons, which could suggest a false positive in some cases. However, many of the tests involve the same models or subsets, meaning the p-values are not independent. As a result, the observed number of significant p-values (¡ 0.04, in fact, ¡ 0.01 for many cases) is less indicative of false positives than it would be under the assumption of independent tests. This reduces the likelihood of false positives in our findings.
Appendix D Human Evaluation
Human performance is lower than expected due to several factors: (1) Annotation was conducted under time constraints, limiting thorough reasoning. (2) Annotators may lack relevant background knowledge—some were non-native English/French speakers, potentially affecting speech or lyric comprehension. (3) Annotators held only associate degrees, which may impact annotation quality in complex reasoning tasks such as music expertise.